# Graham’s rearrangement for dihedral groups

### v1

@@ 1. $(i)$
    for each proper non-empty interval $I=[i,j]\subseteq[1,u]$, we have that
@@ 2. $(ii)$
    $id\not\in\operatorname{IS}(\mathbf{x})\cdot\tau_{u}\cdots\tau_{j}\cdot\left(\left(\prod_{\leq K}(T_{j})\right)^{-1}\cup\prod_{\leq K}(T_{j-1})\right)$
@@ 3. $(iii)$
    the ordering $\tau_{1},\ldots,\tau_{u},\mathbf{x}$ is a sequencing
@@
Thanks to Lemma 5.1 and Lemma 5.1 of [4] we have that $\mathbb{P}(\prod_{i\in I}\tau_{i}=x)\leq e^{-\Omega(R)}$ for any proper nonempty subinterval $I\subset[1,u]$ and any element $x\in Dih_{p}$. Observe that $(iii)$ holds whenever $(i)$ and $(ii)$ hold since we have assumed that $\delta,\mathbf{x}$ is a sequencing. Then since $|T_{j}|=\Theta(R)$ and for $K=o(R)$ we have $|\prod_{\leq K}(T_{j})|=e^{o(R)}$, it is clear that $(i)$ fails for a subinterval $I$ with probability at most $e^{-\Omega(R)}$. Hence, by the union bound over all the proper nonempty subintervals of $[1,u]$, it can be seen that the condition $(i)$ fails with probability $o(1)$ as $p\to\infty$. The proof of $(ii)$ can be easily derived by noticing that $|\operatorname{IS}(\mathbf{x})|\leq e^{c(\log p)^{1/4}}$. ∎
@@
Let $T_{1}$ and $T_{u}$ be the two subsets chosen randomly as described previously. Then with positive probability we have that for $h=1,u$
$\left|\prod_{=j}(T_{h})\cap\left(\delta\cdot\operatorname{IS}(\mathbf{x})\right)\right|\leq 4K\frac{\binom{|T_{j}^{o}|}{\lceil j/2\rceil}\binom{|T_{j}^{e}|}{\lfloor j/2\rfloor}}{\binom{|D_{j}^{o}|}{\lceil j/2\rceil}\binom{|D_{j}^{e}|}{\lfloor j/2\rfloor}}\left|\prod_{=j}(D_{h})\cap\left(\delta\cdot\operatorname{IS}(\mathbf{x})\right)\right|\,,$ (8)
if $T_{h}$ is a dissociated set that is constantly-one in its $\mathbb{F}_{2}$ components, otherwise
$\left|\prod_{=j}(T_{h})\cap\left(\delta\cdot\operatorname{IS}(\mathbf{x})\right)\right|\leq 4K\frac{\binom{|T_{j}|}{j}}{\binom{|D_{j}|}{j}}\left|\prod_{=j}(D_{h})\cap\left(\delta\cdot\operatorname{IS}(\mathbf{x})\right)\right|\,.$ (9)
@@
Here $T_{1}$ is chosen uniformly at random from all the subsets of $D_{1}$ with size equal to $|D_{1}|/4$ if $D_{1}$ is a dissociated set whose $\mathbb{F}_{2}$ components are constantly zero. Otherwise, $T_{1}$ is the union of two sets $T_{1}^{o}$ and $T_{1}^{e}$ where each of these sets is chosen uniformly at random from all the subsets of $D_{1}^{o}$ and $D_{1}^{e}$, respectively, with size equal to $|D_{1}|/8$. One can see that the expected value of the lhs of Inequalities (8) and (9) is exactly the rhs of those equations without the factor $4K$. Therefore, by Markov’s inequality, we see that bounds (8) and (9) fail with probability at most $1/4K$. The same argument applies also for $T_{u}$. Finally, we conclude the proof of this lemma by taking the union bound over all $j\in[1,K]$ and $h=1,u$. ∎
@@
$Y_{j}:=\left(\prod_{=j}(D_{s})\right)^{-1}\cup\left(\delta^{-1}\prod_{=j}(D_{1})\right)$ (10)
for each $1\leq j\leq K$ and we apply Proposition 4.1. This gives us an ordering of $\mathbf{x}$ such that $\delta,\mathbf{x}$ is a sequencing and such that Inequality (4) holds. Then using Lemmas 5.2 and 5.3, we obtain dissociated sets $T_{1},\ldots,T_{u}$ from $D_{1},\ldots,D_{s}$ such that $A=P\cup N\cup S\cup\left(\cup_{i=1}^{u}T_{i}\right)$ satisfy the conditions imposed in these lemmas. Then fixing such a choice for $T_{1},\ldots,T_{u}$, we will construct a sequencing of $A$ of the form $\mathbf{t}_{1},\ldots,\mathbf{t}_{u},\mathbf{x}$, where the $\mathbf{t}_{i}$s are random orderings chosen with a suitable distribution described later. Now, we take care of the remaining Type I intervals left open by Lemmas 5.2 and 5.3. Following the procedure of [4], we pick the orderings of each $\mathbf{t}_{i}$ uniformly at random on a restricted subset of the possible orderings.
We say that an ordering $t_{1},t_{2},\ldots,t_{|T_{i}|}$ of $T_{i}$ is acceptable if
$t_{1}\cdot t_{2}\cdots t_{k}\not\in\delta\cdot\operatorname{IS}(\mathbf{x})\,,$
for every $1\leq k\leq K$. We then state the following lemma, which is an adaptation of Lemma 6.1 of [4].
@@
If $T_{1}$ is a dissociated set that is constantly-zero in its $\mathbb{F}_{2}$ components, pick an ordering of $T_{1}$ uniformly at random from the orderings of $T_{1}$; otherwise pick the elements in odd positions of such an ordering uniformly at random from the orderings of $T_{1}^{o}$ and the elements in even positions uniformly at random from the orderings of $T_{1}^{e}$. Pick a random ordering of $T_{u}$ by the same process. Then these two orderings are acceptable with positive probability.
@@
We prove the statement only for $T_{1}$ since the proof is analogous for $T_{u}$. Let $\mathbf{t}_{1}=t_{1},\ldots,t_{|T_{1}|}$ be a random ordering of $T_{1}$. If $T_{1}$ is a dissociated set that is constantly-zero in its $\mathbb{F}_{2}$ components then $\mathbf{t}_{1}$ is chosen uniformly at random from the orderings of $T_{1}$; otherwise $t_{1},t_{3},\ldots$ are chosen uniformly at random from the orderings of $T_{1}^{o}$ and $t_{2},t_{4},\ldots$ from the orderings of $T_{1}^{e}$. We want to prove that
$\mathbb{P}\left(t_{1}\cdots t_{k}\in\delta\cdot\operatorname{IS}(\mathbf{x})\right)\leq\frac{1}{100K}\,,$
for each $1\leq k\leq K$. For a fixed value of $k$, since $T_{1}$ is a dissociated set, Lemma 3.6 tells us that the products $t_{1}\cdots t_{k}$ are uniformly distributed over the set $\prod_{=k}(T_{1})$. Given the definition of the sets $Y_{j}$s in (10), we see that $|Y_{j}|=2\left|\prod_{=j}(D_{1})\right|$ since $|D_{1}|=|D_{s}|$ and $D_{1}\cup D_{s}\cup\{\delta\}$ is dissociated by condition $(v)$ of the structure theorem. Then by Proposition 4.1, taking $L=\lfloor|Y_{k}|^{1/2}\rfloor$ and $K=c_{2}R^{1/3}$, we have that
$\left|\prod_{=k}(D_{1})\cap\left(\delta\cdot\operatorname{IS}(\mathbf{x})\right)\right|=O\left(|Y_{k}|^{1/2}+\sum_{i=1}^{k-1}|Y_{i}|\right)=O\left(\left|\prod_{=k}(D_{1})\right|\cdot\frac{K}{|D_{1}|}\right)=O\left(\left|\prod_{=k}(D_{1})\right|\cdot c_{2}^{3}K^{-2}\right)\,.$ (11)
Since the rhs of Equation (9) is an upper bound for the rhs of Equation (8), by Lemma 5.3 and Equation (11) we obtain
$\left|\prod_{=k}(T_{1})\cap\left(\delta\cdot\operatorname{IS}(\mathbf{x})\right)\right|=O\left(4K\left|\prod_{=k}(T_{1})\right|c_{2}^{3}K^{-2}\right)=O\left(\left|\prod_{=k}(T_{1})\right|c_{2}^{3}K^{-1}\right)\,.$
@@
$\mathbb{P}\left(t_{1}\cdots t_{k}\in\delta\cdot\operatorname{IS}(\mathbf{x})\right)=O\left(c_{2}^{3}K^{-1}\right)\,.$
Therefore, for a sufficiently small constant $c_{2}$, we have that $\mathbb{P}\left(t_{1}\cdots t_{k}\in\delta\cdot\operatorname{IS}(\mathbf{x})\right)\leq 1/(100K)$. The theorem follows by the union bound over $1\leq k\leq K$ since
$\mathbb{P}\left(\cup_{k=1}^{K}t_{1}\cdots t_{k}\in\delta\cdot\operatorname{IS}(\mathbf{x})\right)\leq\sum_{k=1}^{K}\mathbb{P}\left(t_{1}\cdots t_{k}\in\delta\cdot\operatorname{IS}(\mathbf{x})\right)\leq\frac{1}{100}\,.$
Following [4], we choose $\mathbf{t}_{1},\mathbf{t}_{u}$ independently such that $\mathbf{t}_{1},\overline{\mathbf{t}_{u}}$ are acceptable orderings (here $\overline{\mathbf{t}_{u}}=t_{|T_{1}|},\ldots,t_{1}$ denotes the reverse of $\mathbf{t}_{u}$). This choice takes care of Type I intervals with one endpoint in $\mathbf{x}$ and the other one in the first $K$ elements of $\mathbf{t}_{1}$ or in the last $K$ elements of $\mathbf{t}_{u}$. Now, we choose the pair of orderings $\mathbf{t}_{2j},\mathbf{t}_{2j+1}$ for $j\in[1,u/2-1]$ as follows: Given a partial ordering $t_{1},\ldots,t_{k}$ of $T_{2j}$ where if $T_{2j}$ is a dissociated set with constantly-one $\mathbb{F}_{2}$ components then $t_{1},t_{3},\ldots$ is a partial ordering of $T_{2j}^{o}$ and $t_{2},t_{4},\ldots$ is a partial ordering of $T_{2j}^{e}$, and given a partial ordering $t^{\prime}_{1},\ldots,t^{\prime}_{\ell}$ of $T_{2j+1}$ by the same process, we say that the pair of partial orderings is permissible if $t_{1}\cdots t_{i}\cdot t^{\prime}_{1}\cdots t^{\prime}_{s}\neq id$ for all $(i,s)$. We let $t_{1},t_{2},\ldots,t_{K}$ and $t^{\prime}_{1},t^{\prime}_{2},\ldots,t^{\prime}_{K}$ be a uniformly random permissible pair of orderings, then we choose $\mathbf{t}_{2j}$ to be a uniformly random ordering of $T_{2j}$ if $T_{2j}$ is a dissociated set with constantly-zero $\mathbb{F}_{2}$ components, otherwise the elements in odd positions are picked uniformly at random from $T_{2j}^{o}$ and the elements in even positions from the orderings of $T_{2j}^{e}$ conditional on $\overline{\mathbf{t}_{2j}}$ starting with $t_{1},\ldots,t_{K}$ and $\mathbf{t}_{2j+1}$ is picked with the same process conditional on $\mathbf{t}_{2j+1}$ starting with $t^{\prime}_{1},\ldots,t^{\prime}_{K}$. These choices are made independently for each $j$ and from the choices of $\mathbf{t}_{1},\mathbf{t}_{u}$.
As done in [4], we can see that also in our case, for any $11\leq j\leq u/2-$ 1 and any orderings of $\mathbf{t}_{2j}$ and $\mathbf{t}_{2j+1}$, we have that
@@
$\mathbb{P}\left(\overline{\mathbf{t}_{2j}}=t_{1},\ldots,t_{|T_{2j}|}\mid\mathbf{t}_{2j+1}=t^{\prime}_{1},\ldots,t^{\prime}_{|T_{2j+1}|}\right)=O\left(\frac{1}{|T_{2j}|!}\right)\,.$ (12)
@@
Now, if we choose randomly the orderings $\mathbf{t}_{1},\mathbf{t}_{2},\ldots,\mathbf{t}_{u}$ as described above, proceeding as in Lemma 6.3 of [4] we have with high probability that the following ordering,
$\mathbf{t}_{1},\mathbf{t}_{2},\ldots,\mathbf{t}_{u},\mathbf{x}\,,$
is a sequencing whenever $|\mathbf{x}|\leq e^{c(\log p)^{1/4}}$ and $p$ is sufficiently large. More precisely, we can state the following theorem:
@@
Let $p$ be large enough and let $c>0$ and $1\leq s\leq e^{c(\log p)^{1/4}}$. Let $D_{1},D_{2},\ldots,D_{s}\subseteq Dih_{p}$ be dissociated sets, each of size $\Theta(R)$ a multiple of eight.
Defining $\delta$ as in Theorem 3.5(iv), we assume that $D_{1},D_{s}$, and $\delta$ satisfy condition $(v)$ of that theorem. Now we consider a sequence $\mathbf{x}$ over $Dih_{p}$ of length at most $e^{c(\log p)^{1/4}}$ such that $\delta,\mathbf{x}$ is a sequencing satisfying Condition (4) with respect to the sets $Y_{j}$s defined in Equation (10).
Then it is possible to choose the sequence $T_{1},\ldots,T_{u}$ of dissociated sets and the orderings $\mathbf{t}_{1},\dots,\mathbf{t}_{u}$ as described above so that each $\mathbf{t}_{i}$ is a sequencing and so is
$\mathbf{t}_{1},\ldots,\mathbf{t}_{u},\mathbf{x}.$
This easily implies Theorem 1.3. For reader convenience, we briefly recall below the scheme of the procedure.
Given a set $A\subseteq Dih_{p}\setminus\{id\}$ we first note that, if $A\subseteq\{(x,0):x\in\mathbb{F}_{p}\}$ (which is isomorphic to $\mathbb{F}_{p}$), then the statement of Theorem 1.3 follows from that of Theorem 1.2. Then, we assume, up to automorphism of $Dih_{p}$, that
@@
where the sets $D_{j}$s and $E$ are defined according to Theorem 3.5. If $A=E$ (i.e. $s=0$) Theorem 1.3 immediately follows from Proposition 4.1. Otherwise, we have $s>0$ and, up to changing all the signs of the $\mathbb{F}_{p}$ components, we can assume that $z_{0}>0$ for the $\delta=(z_{0},0)$ defined in Theorem 3.5(iv). Now we apply 4.1 to obtain an ordering $\mathbf{x}$ of $E$ such that $\delta,\mathbf{x}$ is a sequencing satisfying Condition (4) with respect to the sets $Y_{j}$s mentioned above. Now Theorem 1.3 follows since, if we assume that $|A|\leqslant e^{c(\log p)^{1/4}}$, Theorem 5.5 shows the existence of a sequencing of $A$.

### v2

@@ 1. $(i)$
    for each proper nonempty interval $I=[i,j]\subseteq[1,u]$, we have that
@@ 2. $(ii)$
    $id\not\in\operatorname{IS}(\mathbf{x}_{1})\cdot\tau_{1}\cdots\tau_{j}\cdot\left(\left(\prod_{\leq K}(T_{j})\right)^{-1}\cup\prod_{\leq K}(T_{j+1})\right)$
    $id\not\in\operatorname{IS}(\mathbf{x}_{2})\cdot\tau_{u}\cdots\tau_{j}\cdot\left(\left(\prod_{\leq K}(T_{j})\right)^{-1}\cup\prod_{\leq K}(T_{j-1})\right)$
@@ 3. $(iii)$
    the ordering $\mathbf{x}_{1},\tau_{1},\ldots,\tau_{u},\mathbf{x}_{2}$ is a sequencing
@@
Thanks to Lemma 4.4 and Lemma 5.1 of [4], we have that $\mathbb{P}(\prod_{i\in I}\tau_{i}=x)\leq e^{-\Omega(R)}$ for any proper nonempty subinterval $I\subset[1,u]$ and any element $x\in\mathbb{Z}_{p}\rtimes_{\varphi}H$. Observe that $(iii)$ holds whenever $(i)$ and $(ii)$ hold since we have assumed that $\mathbf{x}_{1},\delta,\mathbf{x}_{2}$ is a sequencing. Then since $|T_{j}|\asymp R$ and for $K=o(R)$ we have $|\prod_{\leq K}(T_{j})|=e^{o(R)}$, it is clear that $(i)$ fails for a subinterval $I$ with probability at most $e^{-\Omega(R)}$. Hence, by the union bound over all the proper nonempty subintervals of $[1,u]$, it can be seen that the condition $(i)$ fails with probability $o(1)$ as $p\to\infty$. The proof of $(ii)$ can be easily derived by noticing that $|\operatorname{IS}(\mathbf{x}_{1})|,|\operatorname{IS}(\mathbf{x}_{2})|\leq e^{c(\log p)^{1/4}}$. ∎
@@
Let $T_{1}$ and $T_{u}$ be the two subsets chosen randomly as described previously. Then with positive probability we have for $h=1,u$, that
(15) $\left|\prod_{=j}(T_{h})\cap\left(-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\right|\leq 4K\frac{\binom{|T_{j}^{o}|}{\lceil j/2\rceil}\binom{|T_{j}^{e}|}{\lfloor j/2\rfloor}}{\binom{|D_{j}^{o}|}{\lceil j/2\rceil}\binom{|D_{j}^{e}|}{\lfloor j/2\rfloor}}\left|\prod_{=j}(D_{h})\cap\left(-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\right|\,$
if $T_{h}$ is a dissociated set whose $H$ components are in $H_{1}$. Otherwise
(16) $\left|\prod_{=j}(T_{h})\cap\left(-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\right|\leq 4K\frac{\binom{|T_{j}|}{j}}{\binom{|D_{j}|}{j}}\left|\prod_{=j}(D_{h})\cap\left(-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\right|\,.$
@@
Here $T_{1}$ is chosen uniformly at random from all the subsets of $D_{1}$ with size equal to $|D_{1}|/4$ if $D_{1}$ is a dissociated set whose $H$ components are in $H_{1}$. Otherwise, $T_{1}$ is the union of two sets $T_{1}^{o}$ and $T_{1}^{e}$ where each of these sets is chosen uniformly at random from all the subsets of $D_{1}^{o}$ and $D_{1}^{e}$, respectively, with size equal to $|D_{1}|/8$. One can see that the expected value of the left-hand side of inequalities (15) and (16) is exactly the right-hand side of those equations without the factor $4K$. Therefore, by Markov’s inequality, we see that bounds (15) and (16) fail with probability at most $1/4K$. The same argument applies also for $T_{u}$. Finally, we conclude the proof of this lemma by taking the union bound over all $j\in[1,K]$ and $h=1,u$. ∎
@@
(17) $Y_{j}:=\left(\prod_{=j}(D_{s})\right)^{-1}\cup\left(\delta^{-1}\prod_{=j}(D_{1})\right)$
(18) $Y^{\prime}_{j}:=\left(\prod_{=j}(D_{1})\right)^{-1}\cup\left(\delta^{-1}\prod_{=j}(D_{s})\right)$
for each $1\leq j\leq K$ and we apply Proposition 4.2. This gives us an ordering of $\mathbf{x}$ such that $\delta,\mathbf{x}$ is a sequencing, and inequality (9) holds. Then, using Lemmas 4.5 and 4.6, we obtain dissociated sets $T_{1},\ldots,T_{u}$ from $D_{1},\ldots,D_{s}$ such that $A=P\cup N\cup Z\cup S\cup\left(\cup_{i=1}^{u}T_{i}\right)$ satisfy the conditions imposed in these lemmas. Fixing such a choice for $T_{1},\ldots,T_{u}$, we will construct a sequencing of $A$ of the form $\mathbf{x}_{1},\mathbf{t}_{1},\ldots,\mathbf{t}_{u},\mathbf{x}_{2}$, where the $\mathbf{t}_{i}$s are random orderings chosen with a suitable distribution described later. Now, we take care of the remaining Type I intervals left open by Lemmas 4.5 and 4.6. Following the procedure of [4], we pick the orders of each $\mathbf{t}_{i}$ uniformly at random on a restricted subset of the possible orderings.
We say that an ordering $t_{1},t_{2},\ldots,t_{|T_{1}|}$ of $T_{1}$ is acceptable if
$t_{1}\cdot t_{2}\cdots t_{k}\not\in-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\,$
for every $1\leq k\leq K$, and we say that an ordering $t_{1},t_{2},\ldots,t_{|T_{u}|}$ of $T_{u}$ is acceptable if
$t_{1}\cdot t_{2}\cdots t_{k}\not\in-\operatorname{IS}(\mathbf{x}_{2})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{1})\right)\,$
for every $1\leq k\leq K$. We can now state the following lemma, which is an adaptation of Lemma 6.1 of [4].
@@
If $T_{1}$ is a dissociated set whose $H$ components are in $H_{1}$, pick an ordering of $T_{1}$ uniformly at random from the orderings of $T_{1}$; otherwise pick the elements in odd positions of such an ordering uniformly at random from the orderings of $T_{1}^{o}$ and the elements in even positions uniformly at random from the orderings of $T_{1}^{e}$. Pick a random ordering of $T_{u}$ by the same process. Then these two orderings are acceptable with positive probability.
@@
We prove the statement only for $T_{1}$ since the proof is analogous for $T_{u}$. Let $\mathbf{t}_{1}=t_{1},\ldots,t_{|T_{1}|}$ be a random ordering of $T_{1}$. If $T_{1}$ is a dissociated set whose $H$ components are in $H_{1}$ then $\mathbf{t}_{1}$ is chosen uniformly at random from the orderings of $T_{1}$; otherwise $t_{1},t_{3},\ldots$ are chosen uniformly at random from the orderings of $T_{1}^{o}$ and $t_{2},t_{4},\ldots$ from the orderings of $T_{1}^{e}$. We want to prove that
$\mathbb{P}\left(t_{1}\cdots t_{k}\in-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\leq\frac{1}{100K}\,$
for each $1\leq k\leq K$. For a fixed value of $k$, since $T_{1}$ is a dissociated set, Lemma 3.6 tells us that the products $t_{1}\cdots t_{k}$ are uniformly distributed over the set $\prod_{=k}(T_{1})$. Given the definition of the sets $Y_{j}$s in (17), we see that $|Y_{j}|=2\left|\prod_{=j}(D_{1})\right|$ since $|D_{1}|=|D_{s}|$ and $D_{1}\cup D_{s}\cup\{\delta\}$ is dissociated by condition $(v)$ of the Structure Theorem. Then, by Proposition 4.2, taking $L=\lfloor|H|^{1/2}|Y_{k}|^{1/2}\rfloor$ and $K=c_{2}R^{1/3}$, we have that
$\left|\prod_{=k}(D_{1})\cap\left(-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\right|$ $=O\left(|H|^{1/2}|Y_{k}|^{1/2}+|H|\sum_{i=1}^{k-1}|Y_{i}|+|H|\right)$
$=O\left(\left|\prod_{=k}(D_{1})\right|\cdot\frac{K}{|D_{1}|}\cdot|H|\right)$
(19) $=O\left(\left|\prod_{=k}(D_{1})\right|\cdot|H|\cdot c_{2}^{3}K^{-2}\right)\,.$
Since the right-hand side of Equation (16) is an upper bound for the right-hand side of Equation (15), by Lemma 4.6 and Equation (4) we obtain
$\left|\prod_{=k}(T_{1})\cap\left(-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\right|=O\left(4K\left|\prod_{=k}(T_{1})\right|\cdot|H|\cdot c_{2}^{3}K^{-2}\right)=O\left(\left|\prod_{=k}(T_{1})\right|\cdot|H|\cdot c_{2}^{3}K^{-1}\right)\,.$
@@
$\mathbb{P}\left(t_{1}\cdots t_{k}\in-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)=O\left(|H|\cdot c_{2}^{3}K^{-1}\right)\,.$
Therefore, for a sufficiently small real $c_{2}$ that depends on $|H|$, we have that $\mathbb{P}\left(t_{1}\cdots t_{k}\in\delta\cdot\operatorname{IS}(\mathbf{x})\right)\leq 1/(100K)$. The theorem follows by the union bound over $1\leq k\leq K$ since
$\mathbb{P}\left(\cup_{k=1}^{K}t_{1}\cdots t_{k}\in-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\leq\sum_{k=1}^{K}\mathbb{P}\left(t_{1}\cdots t_{k}\in-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\leq\frac{1}{100}\,.$
Following [4], we choose $\mathbf{t}_{1},\mathbf{t}_{u}$ independently such that $\mathbf{t}_{1},\mathbf{t}_{u}$ are acceptable orderings (here $\overline{\mathbf{t}}_{u}=t_{|T_{1}|},\ldots,t_{1}$ denotes the reverse of $\mathbf{t}_{u}$). This choice takes care of Type I intervals with one endpoint in $\mathbf{x}$ and the other one in the first $K$ elements of $\mathbf{t}_{1}$ or in the last $K$ elements of $\mathbf{t}_{u}$. Now, we choose the pair of orderings $\mathbf{t}_{2j},\mathbf{t}_{2j+1}$ for $j\in[1,u/2-1]$ as follows: given a partial ordering $t_{1},\ldots,t_{k}$ of $T_{2j}$ where if $T_{2j}$ is a dissociated set whose $H$ components are in $H_{1}$ then $t_{1},t_{3},\ldots$ is a partial ordering of $T_{2j}^{o}$ and $t_{2},t_{4},\ldots$ is a partial ordering of $T_{2j}^{e}$, and given a partial ordering $t^{\prime}_{1},\ldots,t^{\prime}_{\ell}$ of $T_{2j+1}$ by the same process, we say that the pair of partial orderings is permissible if $t_{1}\cdots t_{i}\cdot t^{\prime}_{1}\cdots t^{\prime}_{s}\neq id$ for all $(i,s)$. We let $t_{1},t_{2},\ldots,t_{K}$ and $t^{\prime}_{1},t^{\prime}_{2},\ldots,t^{\prime}_{K}$ be a uniformly random permissible pair of orderings, then we choose $\mathbf{t}_{2j}$ to be a uniformly random ordering of $T_{2j}$ if $T_{2j}$ is a dissociated set whose $H$ components are in $H_{0}$. Otherwise, the elements in odd positions are picked uniformly at random from $T_{2j}^{o}$ and the elements in even positions from the orderings of $T_{2j}^{e}$ conditional on $\overline{\mathbf{t}}_{2j}$ starting with $t_{1},\ldots,t_{K}$. Moreover, $\mathbf{t}_{2j+1}$ is picked with the same process conditional on $\mathbf{t}_{2j+1}$ starting with $t^{\prime}_{1},\ldots,t^{\prime}_{K}$. These choices are made independently for each $j$ and from the choices of $\mathbf{t}_{1},\mathbf{t}_{u}$.
As done in [4], we can see that also in our case, for any $1\leq j\leq u/2-1$ and any orderings of $\mathbf{t}_{2j}$ and $\mathbf{t}_{2j+1}$, we have that
@@
(20) $\mathbb{P}\left(\mathbf{t}_{2j}=t_{1},\ldots,t_{|T_{2j}|}\mid\mathbf{t}_{2j+1}=t^{\prime}_{1},\ldots,t^{\prime}_{|T_{2j+1}|}\right)=O\left(\frac{1}{|T_{2j}|!}\right)\,.$
@@
Now, if we choose randomly the orderings $\mathbf{t}_{1},\mathbf{t}_{2},\ldots,\mathbf{t}_{u}$ as described above, proceeding as in Lemma 6.3 of [4], we have with high probability that the ordering
$\mathbf{x}_{1},\mathbf{t}_{1},\mathbf{t}_{2},\ldots,\mathbf{t}_{u},\mathbf{x}_{2}\,$,
is a sequencing whenever $|\mathbf{x}_{1}|,|\mathbf{x}_{2}|\leq e^{c(\log p)^{1/4}}$ and $p$ is sufficiently large. More precisely, we can state the following theorem:
@@
Let $H$ be a finite abelian and strongly sequenceable group. Let $c>0$ and let us consider $\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $\varphi$ satisfies property $(\star)$ and $p$ is large enough depending on $H$. Let $D_{1},D_{2},\ldots,D_{s}\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$ be dissociated sets, each of size $\asymp R$ and a multiple of eight where $1\leq s\leq e^{c(\log p)^{1/4}}$.
Defining $\delta$ as in Theorem 3.5(iv), we assume that $D_{1},D_{s}$, and $\delta$ satisfy condition $(v)$ of that theorem. Now we consider sequences $\mathbf{x}_{1},\mathbf{x}_{2}$ over $\mathbb{Z}_{p}\rtimes_{\varphi}H$ of length at most $e^{c(\log p)^{1/4}}$ such that $\mathbf{x}_{1},\delta,\mathbf{x}_{2}$ is a sequencing satisfying Proposition 4.2 when $\mathbf{x}_{1}=\emptyset$ or Proposition 4.3 when $\mathbf{x}_{1}\neq\emptyset$ with respect to the sets $Y_{j}$s defined in Equation (17) for $\mathbf{x}_{1}=\emptyset$ and with respect to the sets $Y_{j}$s and $Y^{\prime}_{j}$s defined in Equations (17) and (18) for $\mathbf{x}_{1}\neq\emptyset$.
Then it is possible to choose the sequence $T_{1},\ldots,T_{u}$ of dissociated sets and the orderings $\mathbf{t}_{1},\dots,\mathbf{t}_{u}$ so that each $\mathbf{t}_{i}$ is a sequencing and so is
$\mathbf{x}_{1},\mathbf{t}_{1},\ldots,\mathbf{t}_{u},\mathbf{x}_{2}.$
Now, we obtain the main result of our work.
Theorem 1.3.
Let $H$ be a finite abelian and strongly sequenceable group. There exists a $c>0$ such that every subset $A\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H\setminus\{id\}$, where $\varphi:H\to Aut(\mathbb{Z}_{p})$ satisfies $\varphi(h)\in\{id,-id\}$ for each $h\in H$, of size
$|A|\leq e^{c(\log p)^{1/4}},$
has a sequencing.
Proof.
We assume, up to automorphism of $\mathbb{Z}_{p}\rtimes_{\varphi}H$, that
@@
where the sets $D_{j}$s and $E$ are defined according to Theorem 3.5. If $A=E$ (i.e. $s=0$) then Theorem 1.3 follows immediately from Proposition 4.2. Otherwise, we have $s>0$ and, up to changing all the signs of the $\mathbb{Z}_{p}$ components, we can assume that $z_{0}>0$ and $\varphi(a_{0})=id$ for the $\delta=(z_{0},a_{0})$ defined in Theorem 3.5(iv). As done in Equation (6), we can split the set $E$ into four sets $P,N,Z$ and $S$. Now, if $S\neq\emptyset$ then we apply Proposition 4.2 to obtain an ordering $\mathbf{x}$ of $E$ such that $\delta,\mathbf{x}$ is a sequencing satisfying Condition (9) with respect to the sets $Y_{j}$s defined in Equation (17), while for the case $S=\emptyset$, we apply Proposition 4.3 to obtain orderings $\mathbf{z}$ of $Z$, $\mathbf{p}$ of $P$, and $\mathbf{n}$ of $N$ such that $\mathbf{z},\overline{\mathbf{p}},\delta,\mathbf{n}$ is a sequencing satisfying Conditions (10) and (11) with respect to the sets $Y_{j}$s and $Y^{\prime}_{j}$s defined in Equations (17) and (18).
Then, let $d_{1}>0$ and $\bar{p}$ such that Theorem 4.8 holds for $c=d_{1}$ and every $p\geq\bar{p}$. Hence for all $p\geq\bar{p}$, we have that all the subsets $A\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H\setminus\{id\}$ of size $|A|\leq e^{d_{1}(\log p)^{1/4}}$ are sequenceable. For $p<\bar{p}$, we set $d_{2}=\log 2/(\log\bar{p})^{1/4}$ to obtain the upper bound $|A|\leq e^{d_{2}(\log p)^{1/4}}<2$. Clearly, all the subsets $A\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H\setminus\{id\}$ of size at most $1$ are sequenceable. Hence, the statement of the theorem follows by taking $c=\min\{d_{1},d_{2}\}$. ∎
5. Acknowledgements
The first and second authors were partially supported by INdAM–GNSAGA. The authors would also like to thank Noah Kravitz and Benjamin Bedert for useful discussions on this topic.